Production Management:

Given the following table, forecast for the production of shirt using least square method for 2019, 2020, and 2021.Show your soluction.

YEAR PRODUCTION
2012 343.5 PIECES
2013 381.2 PIECES
2014 437.5 PIECES
2015 323.6 PIECES
2016 231.4 PIECES
2017 443.5 PIECES
2018 525.2 PIECES
Answer #1

Here response y = Production

and x = (year - origin) / Interval

year 2015 is the origin as it is in the middle

Constructing the table again with new column x we get,

Year Production (y) x
2012 343.5 PIECES -3
2013 381.2 PIECES -2
2014 437.5 PIECES -1
2015 323.6 PIECES 0
2016 231.4 PIECES 1
2017 443.5 PIECES 2
2018 525.2 PIECES 3

Now creating the linear regression line between x and y we get,

Sry Σί= i r,yi _ n.ry 463.6-7*0*383.7 28 702 = 16.557

and,

30 = у-31-r = 383.7-16.557 * 0 383.7

So the regression line becomes, yi = 383.7 + 16.557 xi

Now for year 2019, 2020 and 2021, x will be 4, 5 and 6 respectively.

Now for year 2019 (x = 4) the forecasted value of Production (y) will be,

y = 383.7 + 16.557 * 4 = 449.928

Now for year 2020 (x = 5) the forecasted value of Production (y) will be,

y = 383.7 + 16.557 * 5 = 466.485

Now for year 2021 (x = 6) the forecasted value of Production (y) will be,

y = 383.7 + 16.557 * 6 = 483.042

**If your answer does not match please comment.